OpenStudy (anonymous):

ALGEBRA 2 HELP PLEASE !!! I don't understand the process I need to do to figure this out ! Can someone please walk me through it ?!?! A cannery processed 1,445 pounds of strawberries in 4.5 hours. The cannery processed 2,320 pounds in 7 hours. Write a linear equation to model the relationship between the weight of strawberries S and time T. How many pounds of strawberries can be processed in 11 hours ?

OpenStudy (mathstudent55):

If you are given two points, such as (3, 5) and (6, 10), do you know how to write the equation of the line that passes through those two points?

OpenStudy (anonymous):

Isn't that when you do \[something = \frac{ y2 - y1 }{ x2-x1 }\] @mathstudent55

OpenStudy (mathstudent55):

Right, you do that to find the slope, m. Then you use one of the points to find the equation using the formula: \( y - y_1 = m(x - x_1)\)

OpenStudy (mathstudent55):

With the example I gave you, you'd do: \(m = \dfrac{10 - 5}{6 - 3} = \dfrac{5}{3}\) Using (3, 5), \(y - 5 = \dfrac{5}{3}(x - 3) \) then you can simplify this to make it look like: 3y - 15 = 5x - 15 3y = 5x \(y = \dfrac{5}{3}x \)

OpenStudy (mathstudent55):

Now back to your problem. Take the given information and use it as two points of a line: (4.5, 1445) and (7, 2320) Those are the two points. Now find the slope and write the equation of the line that passes through them.